# A Graph Neural Network Approach to Dispersed Systems

### v1

@@ Abstract
We present a Graph Neural Network (GNN) that accurately simulates a multidisperse suspension of interacting spherical particles. Our machine learning framework is build upon the recent work of Sanchez-Gonzalez et al. ICML, PMLR, 119, 8459–8468 (2020) on graph network simulators, and efficiently learns the intricate dynamics of the interacting particles. Nodes and edges of the GNN correspond to, respectively, the particles with their individual properties/data (e.g., radius, position, velocity) and the pair interactions between the particles (e.g., electrostatics, hydrodynamics). A key contribution of our work is to account for the finite dimensions of the particles and their impact on the system dynamics. We test our GNN against an exemplary case study of a multidisperse mixture of two-dimensional spheres sedimenting under gravity in a liquid and interacting with each other by a Lennard-Jones potential. The present GNN framework offers a fast and accurate method for the theoretical study of complex physical systems such as field-induced behavior of colloidal suspensions and ionic liquids. Our implementation of the GNN is available on GitHub at github.com/rfjd/GNS-DispersedSystems.
@@ I Introduction
Accurate and efficient simulation of suspensions and emulsions (e.g., colloidal dispersions [1]) is a key component in better understanding the underlying physics of such systems, as well as leveraging them for practical applications spanning from engineering to biology such as cargo-loading and targeted drug-delivery by micro/nano-robots [2, 3, 4], field-induced self-assembly [5, 6], and electrophoretic deposition [7, 8, 9]. Such problems are governed by rather complex, and at times poorly understood, physics. As a result, it is often difficult to develop a comprehensive mathematical model without invoking unrealistic assumptions. Furthermore, even the reduced, approximate, models are often impossible to solve accurately. An alternative approach that circumvents these challenges is to use Machine Learning (ML) to develop a data-driven model.
In a recent seminal work, Sanchez-Gonzalez et al. [10] used Graph Neural Networks (GNNs) [11, 12, 13] to simulate the complex physics of fluid dynamics in a wide range of conditions. The corresponding machine learning framework, referred to as a Graph Network Simulator (GNS), treats the fluid medium (along with the immersed solid objects) as an ensemble of point particles interacting with each other. As such, GNS is similar to particle-based methods such as smoothed particle hydrodynamics [14], position based dynamics [15], and material point method [16].
@@
We extend the GNS [10, 17] by accounting for the finite size of the particles in a fluid suspension. Our methodology acts as an accurate, efficient, surrogate to the often time-consuming, conventional, computational methods such as computational fluid dynamics (CFD). We represent the immersed particles as nodes of a graph, along with their important dynamic and intrinsic properties such as position, velocity, and radius. The interaction between the particles is modeled as edges of the graph; i.e, the pair-interaction between the particles is represented as a message-passing process between the nodes of the graph.
We evaluate the accuracy of our GNN in an exemplary case-study of a multidisperse suspension of particles sedimenting under gravity in a box of fluid. A comparison between the ground-truth data and predictions of the trained GNN indicates a great agreement. The trained GNN accurately and efficiently captures the most intricate characteristics of the system behavior such as individual particle trajectories, clusters, collisions, and boundaries. Our implementation of the GNN is publicly available on GitHub at github.com/rfjd/GNS-DispersedSystems.
II Problem Statement & GNN
@@ II.1 Multidisperse System of Interacting Spheres
We consider $N$ spherical particles (or droplets) with finite radii $a_{i},\;i=1,2,\cdots,N$, interacting and moving in a box of liquid (see Fig. 1). Depending on the size and charge of the particles, and the type of the liquid, various interactions such as electrostatics, hydrodynamics, steric, and van der Waals forces can be present. Also, the particles experience a drag force from the liquid medium, and may move due to spatially and temporally uniform or nonuniform external forces such as those from a gravitational field or an imposed electric field.
The training data contains trajectories of all particles with a time-step $\Delta t$. Such training data can be obtained from experiments or computer simulations. The objective is to train a GNN that understands the dynamics of such system and can accurately predict the trajectories of all particles. In case of computer simulations, the idea is that the trained GNN (i.e., our data-driven model) is much faster than the physics-based simulator. As shown in Fig. 1, we denote the position and velocity of particle $i$ with $\boldsymbol{x}_{i}$ and $\boldsymbol{v}_{i}$, respectively. Furthermore, $\boldsymbol{F}_{ij}$ is the force on particle $i$ due to the neighboring particle $j$. Of course, we assume that we do not have any information about the nature of this force; instead we see the outcome of the interaction through the trajectories of the particles. Following Sanchez-Gonzalez et al. [10], we use a first-order finite difference scheme to extract particle velocities $\boldsymbol{v}_{i}$ from positions $\boldsymbol{x}_{i}$.
For a physics-based model, one would need two initial conditions (usually, the current positions and velocities) to predict the next position of the particles. Since we use a first-order finite difference scheme to compute the velocities from positions, this translates to a sequence of two positions of all particles (i.e., $\{\boldsymbol{x}(t-2\Delta t),\boldsymbol{x}(t-\Delta t)\}\to\boldsymbol{x}(t)$). Indeed, this information is sufficient for the GNN to predict the particle trajectories as well. However, longer sequences provide significantly better predictions [10]. Unless stated otherwise, for the results presented here, we have used a sequence of $C=6$ last positions to predict the next position. Note that, therefore, a single full-length computational or experimental positional data with $M$ time-steps results in $(M-6)$ training trajectories (or ‘examples’).
@@ II.2 Node & Edge Features
Node Features. We use the last $C-1$ velocities of each particle $\boldsymbol{v}_{i}$ ($(C-1)\times\mathrm{DIM}$ values, where $\mathrm{DIM}$ denotes the spatial dimension of the system), its distance to boundaries (shown as $\ell_{i}$’s in Fig. 2; $2\times\mathrm{DIM}$ values), and its radius $a_{i}$ as node features. We also normalize the velocities and distances corresponding to each particle by its radius. We acknowledge that a more physically appropriate normalization would be a constant length scale for all particles. But our scaling, along with what we use for the edge features (see next paragraph), helps the GNN to accurately learn the important characteristics of the system (e.g., clustering).
Edge Features. We consider relative displacement, $\boldsymbol{x}_{ij}=\boldsymbol{x}_{i}-\boldsymbol{x}_{j}$, and absolute distance, $\left\|\boldsymbol{x}_{ij}\right\|$, between particles (see Fig. 2; $\mathrm{DIM}+1$ values), last $C-1$ relative velocities $\boldsymbol{v}_{i}-\boldsymbol{v}_{j}$ ($(C-1)\times\mathrm{DIM}$ values), and sum of the particle radii for each particle pair, $a_{i}+a_{j}$, ($1$ value) as the edge features. We normalize all lengths by $a_{i}+a_{j}$.
@@ II.3 GNN Architecture
Our GNN architecture is largely similar to that developed by Sanchez-Gonzalez et al. [10], and in particular its PyTorch implementation by Kumar and Vantassel [17]. An outline of our network architecture is provided in Fig. 3. The node and edge feature vectors ($\boldsymbol{v}_{\mathrm{node}}$ and $\boldsymbol{v}_{\mathrm{edge}}$, respectively) get encoded by two Multilayer Perceptrons (MLPs) into their corresponding embeddings $\boldsymbol{\chi}\in\mathbb{R}^{N\times n_{\mathrm{node}}}$ and $\boldsymbol{e}\in\mathbb{R}^{E\times n_{\mathrm{edge}}}$, respectively. Here, symbols stand for number of nodes, $N$; number of edges, $E$; node embedding size, $n_{\mathrm{node}}$; and edge embedding size, $n_{\mathrm{edge}}$. The node and edge embedding sizes are both $128$ and kept the same throughout the network. In each step of the message passing from a neighbor node $j$ to a target node $i$, we concatenate $\boldsymbol{\chi}_{i}$, $\boldsymbol{\chi}_{j}$, and $\boldsymbol{e}_{ij}$ and pass it through an MLP. Then we aggregate all messages from the neighbors of the target node $i$ (i.e, $j\in\mathcal{N}(i)$) to find the overall message $\tilde{\boldsymbol{\chi}}\in\mathbb{R}^{N\times n_{\mathrm{edge}}}$. Note that aggregating all messages is inspired by the physics of the problem; all neighboring particles exert force on the target particle, and the overall force is the summation of all of those interaction forces. Following prior work, we finally apply a residual connection to update the node and edge embeddings. The message passing is performed for $K$ steps. The obtained node embedding $\{\boldsymbol{\chi}\}^{K}$ is then passed through a final MLP to predict the particle accelerations $\hat{\ddot{\boldsymbol{x}}}\in\mathbb{R}^{N\times\mathrm{DIM}}$.
The objective of the model is to predict the acceleration of the particles, given a sequence of $C$ last positions. The loss function to minimize during training is
@@
where $\ddot{\boldsymbol{x}}$ and $\hat{\ddot{\boldsymbol{x}}}$ are the true and predicted accelerations, and $N$ is the number of particles in example. Following Sanchez-Gonzalez et al. [10], we corrupt the sequence of particle positions with a random walk noise, and use finite difference to compute the accelerations from true positions. For rollout predictions, we use the same finite difference stencil to compute the particle positions form the predicted accelerations.
@@
Figure 3: An outline of our network architecture including the encoding, message passing, and decoding steps.
III Case Study
Here we consider a case study to put our GNN into test: a multidisperse suspension of spherical disks interacting and sedimenting under gravity in a two-dimensional box $x,y\in[-L,L]$. We perform computer simulations to generate the training and test data. We stress that the following case study is not meant to be a physically accurate representation of the system. However, it has the key characteristics to challenge our GNN model, such as particles of different sizes, intricate interactions between the particles, and physical boundaries. Alternatively, one could collect actual experimental data or some simulation data from a black-box multi-physics simulator to train the model.
III.1 Governing equations
The deterministic dynamics of the particles are governed by the Newton’s second law

### v3

# A Graph Neural Network Simulation of Dispersed Systems

@@ Abstract
We present a Graph Neural Network (GNN) that accurately simulates a multidisperse suspension of interacting spherical particles. Our machine learning framework is built upon the recent work of Sanchez-Gonzalez et al. ICML, PMLR, 119, 8459–8468 (2020) on graph network simulators, and efficiently learns the intricate dynamics of the interacting particles. Nodes and edges of the GNN correspond, respectively, to the particles with their individual properties/data (e.g., radius, position, velocity) and the pairwise interactions between the particles (e.g., electrostatics, hydrodynamics). A key contribution of our work is to account for the finite dimensions of the particles and their impact on the system dynamics. We test our GNN against a representative case study of a multidisperse mixture of two-dimensional spheres sedimenting under gravity in a liquid and interacting with each other by a Lennard-Jones potential. The present GNN framework offers a fast and accurate method for the theoretical study of complex physical systems such as field-induced behavior of colloidal suspensions and ionic liquids. Our implementation of the GNN is available on GitHub at github.com/rfjd/GNS-DispersedSystems.
@@ I Introduction
Accurate and efficient simulation of suspensions and emulsions (e.g., colloidal dispersions [1]) is essential for better understanding the underlying physics of these systems, as well as leveraging them for practical applications. These applications span diverse fields, from engineering to biology, including cargo-loading and targeted drug-delivery by micro/nano-robots [2, 3, 4], field-induced self-assembly [5, 6], and electrophoretic deposition [7, 8, 9].
Such problems are governed by complex and often poorly understood physics, making it challenging to develop comprehensive mathematical models without relying on unrealistic assumptions. For instance, several long-standing open questions persist regarding the field-induced behavior of colloidal dispersions. Despite extensive research over the past two decades, theoretical and computational predictions remain ‘qualitatively’ at odds with the experimental observations. For example, computational studies based on current theoretical models fail to predict even the direction of flow in microfluidic systems (AC electroosmosis pumps), let alone its magnitude, highlighting fundamental inaccuracies in the underlying theory [10, 11, 12, 13]. Similarly, available theories in electrokinetics fail to predict some of the experimentally established behaviors of charged colloids near electrodes, including the electrolyte-dependent aggregation or separation of colloids [14, 15] and the extreme levitation of colloids against gravity [16, 17, 15, 18]. To further compound matters, there is significant uncertainty regarding the validity of different modeling frameworks (e.g., continuum, Brownian Dynamics, or Molecular Dynamics), and no consensus exists among researchers on which framework is most suitable for various problems.
Moreover, even if there is a clear understanding of the underlying physics and a comprehensive mathematical model exists, one must accurately and efficiently solve the governing equations to simulate the problem and make predictions. A key challenge in this step is the computational cost, which depends on several factors, including the numerical algorithm (e.g., whether spectral or finite difference/element methods are used, the order of the finite difference/element stencils, and the choice of integration methods), the problem’s complexity (e.g., whether the system involves multiphase flows, turbulence, compressible fluids, or specific types of interactions), the available computational resources, and the implementation details (e.g., programming language and use of parallel computing). Accurate simulation of multiphysics problems can take anywhere from hours to weeks or even months to perform. See a number of specific examples in water/oil emulsion [19], turbulent multiphase flows [20], colloidal dispersions [18], and fluidized beds [21, 22, 23]. For the latter, as communicated by the corresponding authors, the computation time can range from several hours [21] to weeks [22] and even a few months [23], depending on factors such as the number of particles, simulation time, numerical algorithm, and computational resources. Note further that in some cases, the computational cost renders any realistic simulation of the problem virtually impossible. See the concluding remarks in Bukosky et al. [18] for an example.
An alternative approach that circumvents these challenges is to use Machine Learning (ML) to develop data-driven models. The incorporation of ML techniques has transformed the simulation and modeling of multiphysics systems, leading to significant improvements in computational efficiency, scalability, and generalization. These advancements can be broadly classified into two categories. The first category involves hybrid methods that combine ML with traditional Computational Fluid Dynamics (CFD) frameworks. A prominent example is the work of Kochkov et al. [24], where ML algorithms were integrated into traditional CFD solvers to replace components that are sensitive to coarse resolutions. This approach achieves 4040404080808080 times computational speed-ups with grids up to $10$ times coarser, while maintaining high accuracy, stability, and generalization to unseen flow conditions.
The second category consists of purely data-driven approaches. In a recent seminal work, Sanchez-Gonzalez et al. [25] used Graph Neural Networks (GNNs) [26] to simulate the complex physics of fluid dynamics under a wide range of conditions. The corresponding ML framework, referred to as a Graph Network Simulator (GNS), treats the fluid medium (along with the immersed solid objects) as an ensemble of point particles interacting with each other. In this regard, GNS is similar to particle-based methods such as smoothed particle hydrodynamics [27], position based dynamics [28], and material point method [29], which also model materials as ensembles of interacting point particles. This framework has since been extended in various directions. For instance, Sergeev et al. [30] proposed a hierarchical GNN with self-attention pooling, which further improves scalability and accuracy for complex fluid simulations. Bhattoo et al. [31] introduced the idea of embedding Lagrangian mechanics directly into the graph structure, enabling interpretable modeling of kinetic and potential energy alongside dissipative forces. When long-range, many-body interactions are crucial, Ma et al. [32] introduced the idea of higher-order graph connectivity for hydrodynamic interactions in Stokes suspensions. Also, Martinkus et al. [33] developed a hierarchical GNN capable of reducing particle simulation complexity to linear time and space, enabling efficient large-scale simulations of such systems. On the other hand, when extremely short-range and localized interactions are dominant, Filiatraut et al. [34] proposed a framework utilizing a neighborhood selection process, where each selected neighborhood is assigned its own dedicated GNN, for the study of self-assembly in microscale systems. These innovations highlight the growing impact of hybrid ML and GNN approaches across diverse scales and systems.
The main objective of the present work is to extend the GNS [25, 35] by accounting for the finite size of the particles (droplets) in a fluid suspension (emulsion), as opposed to treating them as point particles. This extension centers on designing an innovative set of node and edge features, along with slight modifications to the network architecture, enabling the GNN to robustly learn the effects of particle size on particle-particle and particle-wall interactions. Our methodology serves as an accurate, efficient, surrogate for the often time-consuming, conventional, computational methods, such as CFD, for the study of dispersed systems. The implications of our work are vast. First, as mentioned above, the behavior of colloidal dispersions under electric fields is poorly understood, leaving a number of pressing open questions that hinder the development of electrokinetic systems for important applications in biology and engineering, such as targeted drug delivery and self-assembly. Second, our approach accommodates particles of varying sizes, allowing for the simulation of multidisperse systems. This capability is crucial because many real-world dispersed systems, such as oil-water emulsions in water purification and essential oil extraction, as well as granulation processes in fluidized beds, are inherently multidisperse. Third, the typical computation time for the simulation of such dispersed systems under practical conditions ranges from weeks to months. Our GNN framework can learn the complex physics of these systems from experimental data (or accurate simulation data if applicable) and accurately simulate their dynamics within seconds.
We represent the immersed particles as nodes of a graph, along with their key dynamic and intrinsic properties such as position, velocity, and radius. The interaction between the particles is modeled as edges of the graph; i.e., the pair-interaction between the particles is represented as a message-passing process between the corresponding nodes of the graph. We evaluate the accuracy of our GNN through a representative case-study of a multidisperse suspension of particles sedimenting under gravity in a box of fluid. A comparison between the ground-truth data and the predictions of the trained GNN shows excellent agreement. The trained GNN accurately and efficiently captures the most intricate characteristics of the system behavior, including individual particle trajectories, clusters, collisions, and boundaries. Our implementation of the GNN is publicly available on GitHub at github.com/rfjd/GNS-DispersedSystems.
The remainder of this article is structured as follows. In Sec. II, we describe the Graph Neural Network (GNN) framework that we have developed for this study, including the network architecture and the specific node and edge features used to model the system dynamics and particle interactions. Sec. III presents a case study of a multidisperse suspension sedimenting under gravity in a box of fluid, where we use a numerical solution to generate the sample training data. In Sec. IV, we evaluate the performance of the GNN in accurately predicting the dynamics of the system and compare its results to the ground-truth data. Finally, in Sec. V, we discuss the implications of our findings, address potential limitations, and provide concluding remarks on the utility and future applications of our approach.
@@
II GNN Framework
@@ II.1 Multidisperse System of Interacting Spheres
We consider $N$ spherical particles (or droplets) with finite radii $a_{i},\;i=1,2,\cdots,N$, interacting and moving in a box of liquid (see Fig. 1). Depending on the size and charge of the particles, and the type of the liquid, various interactions such as electrostatics, hydrodynamics, steric, and van der Waals forces can be present. Also, the particles experience a drag force from the liquid medium, and may move due to spatially and temporally uniform or nonuniform external forces, such as those from a gravitational field or an imposed electric field.
The training data contains trajectories of all particles with a time-step $\Delta t$. Such training data can be obtained from experiments or computer simulations. The objective is to train a GNN to learn the dynamics of such a system and accurately predict the trajectories of all particles. In the case of computer simulations, the idea is that the trained GNN (i.e., our data-driven model) is much faster than the physics-based simulator. If there are no accurate computer simulations for a particularly complex problem (due to a knowledge gap in theory or high computational cost), experimental data can be used. As shown in Fig. 1, we denote the position and velocity of particle $i$ with $\boldsymbol{x}_{i}$ and $\boldsymbol{v}_{i}$, respectively. Furthermore, $\boldsymbol{F}_{ij}$ is the force on particle $i$ due to the neighboring particle $j$. We assume that we do not have any information about the nature of this force; instead, we see the outcome of the interaction through the trajectories of the particles. Following Sanchez-Gonzalez et al. [25], we use a first-order finite difference scheme to extract particle velocities $\boldsymbol{v}_{i}$ from positions $\boldsymbol{x}_{i}$.
For a physics-based model, predicting the next position of the particles requires two initial conditions, typically the current positions and velocities. Using a first-order finite difference scheme to compute velocities from positions, this translates to a sequence of two positions of all particles (i.e., $\{\boldsymbol{x}(t-2\Delta t),\boldsymbol{x}(t-\Delta t)\}\to\boldsymbol{x}(t)$). Indeed, this information is sufficient for the GNN to predict the particle trajectories as well. However, using longer sequences has been shown to significantly improve the prediction accuracy [25]. Unless stated otherwise, for the results presented here, we have used a sequence of $C=6$ most recent positions to predict the next position. Consequently, a single full-length computational or experimental positional dataset with $M$ time steps results in $(M-6)$ training trajectories (or ‘examples’).
@@ II.2 Node & Edge Features
Node Features. The node features consist of (i) the last $C-1$ velocities of each particle, $\boldsymbol{v}_{i}$ ($(C-1)\times D$ values, where $D$ denotes the spatial dimension of the system), (ii) the distances to the boundaries ($\ell_{i}$’s in Fig. 2; $2\times D$ values), and (iii) the particle radius $a_{i}$ ($1$ value). These features are selected to characterize each particle’s location relative to the boundaries, and the evolution of its velocity over the last $C-1$ steps. To facilitate learning, we normalize the velocities and distances corresponding to each particle by its radius. We acknowledge that a more physically appropriate normalization would be a constant length scale for all particles. But our scaling approach, combined with the method for defining the edge features (described in the next paragraph), proves effective in enabling the GNN to accurately learn critical system characteristics, such as clustering.
Edge Features. The edge features for a pair of interacting particles $i$ and $j$ include (i) the relative displacement $\boldsymbol{x}_{ij}=\boldsymbol{x}_{i}-\boldsymbol{x}_{j}$ and absolute distance $\left\|\boldsymbol{x}_{ij}\right\|$ (see Fig. 2; $D+1$ values), (ii) the last $C-1$ relative velocities $\boldsymbol{v}_{i}-\boldsymbol{v}_{j}$ ($(C-1)\times D$ values), and (iii) the sum of the particle radii $a_{i}+a_{j}$ ($1$ value). We normalize all velocities and distances by $a_{i}+a_{j}$. These features capture the relative positions of particles and their motion with respect to each other over the last $C-1$ steps. This combination of features ensures that the GNN incorporates the key factors influencing pairwise interactions, such as size asymmetry, proximity, and relative dynamics.
@@ II.3 GNN Architecture
Our GNN architecture is largely similar to that developed by Sanchez-Gonzalez et al. [25], and in particular its PyTorch implementation by Kumar and Vantassel [35]. An overview of the network architecture is provided in Fig. 3. We store the node and edge features as one-dimensional column vectors $\boldsymbol{v}_{\mathrm{node}}\in\mathbb{R}^{(C+1)D+1}$ and $\boldsymbol{v}_{\mathrm{edge}}\in\mathbb{R}^{CD+2}$, respectively. These vectors are encoded by two Multilayer Perceptrons (MLPs) into their corresponding embeddings $\boldsymbol{\chi}\in\mathbb{R}^{N\times n_{\mathrm{node}}}$ and $\boldsymbol{e}\in\mathbb{R}^{E\times n_{\mathrm{edge}}}$, respectively. Here, symbols stand for number of nodes, $N$; number of edges, $E$; node embedding size, $n_{\mathrm{node}}$; and edge embedding size, $n_{\mathrm{edge}}$. The node and edge embedding sizes are set to $128$ and kept the same throughout the network. We then perform a total of $K$ message passing steps as follows. In each step of the message passing from a neighbor node $j$ to a target node $i$, we concatenate $\boldsymbol{\chi}_{i}$, $\boldsymbol{\chi}_{j}$, and $\boldsymbol{e}_{ij}$ and pass it through an MLP. Then we aggregate all messages from the neighbors of the target node $i$ (i.e, $j\in\mathcal{N}(i)$) to find an overall message $\tilde{\boldsymbol{\chi}}\in\mathbb{R}^{N\times n_{\mathrm{edge}}}$. Note that the aggregation of all messages is inspired by the physics of the problem, where each neighboring particle exerts a force on the target particle, and the total force is the sum of these interaction forces. Following prior work, we finally apply a residual connection to update the node and edge embeddings. The final node embedding after $K$ message passing steps, $\{\boldsymbol{\chi}\}^{K}$, is then passed through a final MLP to predict the particle accelerations $\hat{\ddot{\boldsymbol{x}}}\in\mathbb{R}^{N\times D}$.
Figure 3: An outline of our network architecture, including the encoding, message passing, and decoding steps.
During training, the objective of the model is to predict the acceleration of the particles, given a sequence of the last $C$ positions. The loss function to minimize during training is
@@
where $\ddot{\boldsymbol{x}}$ and $\hat{\ddot{\boldsymbol{x}}}$ are the true and predicted accelerations, and $N$ is the number of particles in the example. Following Sanchez-Gonzalez et al. [25], we corrupt the sequence of particle positions with a random walk noise, and use finite difference to compute the accelerations from true positions. For rollout predictions during inference, we use the same finite difference stencil to compute the particle positions from the predicted accelerations.
It is important to note that similar to the prior work by Sanchez-Gonzalez et al. [25], our GNN implementation assumes a time-step of $\Delta t=1$. Consequently, particle accelerations at any step $p$ are computed as $\ddot{\boldsymbol{x}}(p)=\boldsymbol{x}(p)-2\boldsymbol{x}(p-1)+\boldsymbol{x}(p-2)$ (second-order backward finite difference scheme with a unit time-step). Note that this does not cause any issues with the training. When interpreting the network output during inference, the true time-step $\Delta t$ can be incorporated to make predictions in correct units and dimensions.
@@
III Sample Dataset Generation
In this section, we use a toy model to generate synthetic ‘ground truth’ data for training and testing our GNN. The model represents a multidisperse suspension of spherical disks interacting and sedimenting under gravity within a two-dimensional box $x,y\in[-L,L]$. Computer simulations are performed to generate the corresponding training and test datasets. It is important to note that this case study is not intended to provide a physically accurate representation of the system. However, it has the key characteristics to challenge our GNN model, such as particles of different sizes, intricate interactions between the particles, and physical boundaries. Alternatively, one could generate the training and test datasets from actual experimental data or by using a black-box multiphysics simulator like Ansys Fluent.
We describe the motion of the particles using the Newton’s second law: